Problem 10-14 Project Evaluation [LO1]

Dog Up! Franks is looking at a new sausage system with an installed cost of $515,000. This cost will be depreciated straight-line to zero over the project’s five-year life, at the end of which the sausage system can be scrapped for $81,000. The sausage system will save the firm $153,000 per year in pretax operating costs, and the system requires an initial investment in net working capital of $32,000. If the tax rate is 22 percent and the discount rate is 10 percent, what is the NPV of this project? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

NPV
Answer #1
Depreciation per year = 515000 / 5 103000
Annual cashflow = Annual pretax operating cost saving*(1-tax%) + Depreciation*Tax% = 153000*(1-22%) + 103000*22% 142000
After-tax salvage value = 81000*(1-22%) 63180
NPV = 142000*(1-(1+10%)^-5)/10% + 32000/(1+10%)^5 + 63180/(1+10%)^5 - 515000 - 32000 50391.01
If you have any doubt then please ask
Please do rate the answer